# Polynomial stability of non-linearly damped contraction semigroups

### v1

@@ 1. Introduction
We consider semi-linear systems of the form
@@
to be solved subject to the initial condition $x(0)=x_{0}$. Here $A$ is assumed to be the generator of a contraction semigroup $(T(t))_{t\geq 0}$ on a Hilbert space $X$, $x_{0}\in X$ and $B$ is a bounded linear operator from another Hilbert space $U$ into $X$. Furthermore, $\phi\colon U\to U$ is a potentially non-linear map which satisfies $\phi(0)=0$ and is monotone in the sense that
@@
Our results augment earlier studies on asymptotic stability with non-linear damping in [14, 15, 22, 24, 30] by providing rates of decay. They in particular generalise analogous results obtained previously in the linear case [7]. The results in [7] have previously been extended to abstract wave-type equations with non-linear viscous damping in [3, 4, 5, 29]. Rational and generalised decay rates for wave equations with non-linear viscous damping have also been studied in [2, 12, 21], the case of boundary damping in [1, 23, 35, 36, 37]. Exponential stability of wave equations and semigroups with non-linear damping has been investigated for instance in in [13, 19, 25, 34]. Our results differ from existing work in that we establish rational decay rates for general abstract systems with non-linear damping based on a generalised observability property of the linear system. In particular, our class of operators and the type of damping differ from those considered in the most closely related references [3, 4, 5, 29].
A function $x(\cdot)\in C^{1}(\mathbb{R}_{+};X)$ is said to be a classical solution of (1.1) if $x(t)\in D(A)$ for all $t\geq 0$ and (1.1) holds for all $t\geq 0$. A function $x(\cdot)\in C(\mathbb{R}_{+};X)$ is said to be a generalised solution of (1.1) if there exists a sequence $(\smash{x_{n}}(\cdot))_{n\geq 1}$ of classical solutions of (1.1) such that $\sup_{t\in[0,\tau]}\|x(t)-\smash{x_{n}}(t)\|\to 0$ as $n\to\infty$ for all $\tau>0$. Note that if $\phi$ is a linear map then this notion of a classical solution is consistent with the terminology used for linear abstract Cauchy problems, while the notion of a generalised solution describes what would typically be referred to as a mild solution in the linear setting. We recall, moreover, that in the linear case the mild (or generalised) solution is given by the semigroup orbit, and that it is a classical solution precisely when $x_{0}\in D(A)$.
@@
and in particular monotonicity of $\phi$ implies that any classical solution has non-increasing norm. Thus [16, Thm. 11.1.5(b)] shows that (1.1) has a unique classical solution whenever $\phi$ is locally Lipschitz continuous and $x_{0}\in D(A)$. In fact, by monotonicity of $\phi$ any two classical solutions $x_{1}(\cdot),x_{2}(\cdot)$ satisfy
@@
and by [26, Cor. 3.7] the function $\|\dot{x}(\cdot)\|$, too, is non-increasing.
@@ 2. Polynomial stability of abstract semi-linear systems
Suppose that $\phi:U\to U$ is monotone, locally Lipschitz continuous and satisfies $\phi(0)=0$, and that (1.2) holds for some $\delta>0$. Suppose further that there exist $\beta,\tau,c_{\tau}>0$ such that
@@
Then $i\mathbb{R}\subseteq\rho(A_{B})$. Furthermore, all generalised solutions of (1.1) satisfy $\|x(t)\|\to 0$ as $t\to\infty$, and all classical solutions of (1.1) satisfy $\|x(t)\|=O(t^{-1/(2\beta)})$ as $t\to\infty$.
@@
where the implicit constants are independent of $k$. Noting that $x(k\tau)\in D(A)$, we may apply the moment inequality [17, Thm. II.5.34] to obtain
@@
for all sufficiently large $k\geq 1$ by [6, Lem. 1.3.4]. By monotonicity of $\|x(\cdot)\|$ we deduce that
@@
for all sufficiently large $t\geq 0$. Next we observe that, for any $\delta\geq\|B^{*}\|\|x_{0}\|$, the identity function on $U$ defined by $\phi(u)=u$ for all $u\in U$ satisfies our assumptions with $t_{0}=0$ and $L_{\delta}=1$. We deduce that $\|T_{B}(t)(I-A_{B})^{-1}\|=O(t^{-1/(2\beta)})$ as $t\to\infty$, where $(T_{B}(t))_{t\geq 0}$ is the $C_{0}$-semigroup of contractions generated by $A_{B}$. It follows from [9, Thm. 1.1] that $i\mathbb{R}\subseteq\rho(A_{B})$. Moreover, $(T_{B}(t))_{t\geq 0}$ is strongly stable in the sense that $\|T_{B}(t)x\|\to 0$ as $t\to\infty$ for all $x\in X$, by a standard density argument. It then follows from [15, Thm. 2.2] (noting that the result carries over, with the appropriate modifications, to the setting of complex Hilbert spaces, and that the assumptions of compact resolvent and approximate observability can be replaced by strong stability of $(T_{B}(t))_{t\geq 0}$) that whenever $\phi$ satisfies the conditions of our theorem we have $\|x(t)\|\to 0$ as $t\to\infty$ for all generalised solutions of (1.1). Hence the first part of the proof shows that $\|x(t)\|=O(t^{-1/(2\beta)})$ as $t\to\infty$ for all classical solutions of (1.1), which completes the proof of the first part of the result.
Now assume, in addition, that $\|(is-A_{B})^{-1}\|\lesssim 1+|s|^{\alpha}$ for some $\alpha>0$ and all $s\in\mathbb{R}$, and that there exist $\gamma>\alpha/2+1$ and $\kappa,\varepsilon>0$ such that (2.2) holds. If $\alpha\geq 2\beta$ then the result already follows without any of the additional assumptions from what has already been proved, so we may assume that $\alpha<2\beta$. Recall from the first part of the proof that $\|x(t)\|\to 0$ as $t\to\infty$ for all generalised solutions of (1.1), and that $\|x(t)\|=O(t^{-1/(2\beta)})$ as $t\to\infty$ for all classical solutions of (1.1). Note also that by [11, Lem. 2.11(c)] the operator $A_{\kappa}=A-\kappa BB^{\ast}$ satisfies $i\mathbb{R}\subseteq\rho(A_{\kappa})$ and $\|T_{\kappa}(t)A_{\kappa}^{-1}\|=O(t^{-1/\alpha})$ as $t\to\infty$, where $(T_{\kappa}(t))_{t\geq 0}$ is the contraction semigroup generated by $A_{\kappa}$. Let $x(\cdot)$ be a classical solution of (1.1), and let $t_{0}\geq 0$ be such that $\|B^{*}x(t)\|\leq\varepsilon$ for all $t\geq t_{0}$. We have
@@
for all $t\geq t_{0}$. Since $x(t)\in D(A)$ for all $t\geq 0$, we have $\|T_{\kappa}(t-t_{0})x(t_{0})\|=O(t^{-1/\alpha})$ as $t\to\infty$. Moreover, since $\phi$ is locally Lipschitz continuous and $Bx(\cdot)$ is continuously differentiable, we have $\int_{t_{0}}^{t_{\theta}}T_{\kappa}(t_{\theta}-s)By(s)\,\mathrm{d}s\in D(A)$ for all $t\geq t_{0}$ by [8, Cor. 3.1.17]. The second term in (2.3) therefore satisfies
@@
where $y_{\theta}(t)=0$ for $t\in[0,t_{\theta})$ and $y_{\theta}(t)=y(t)$ for $t\geq t_{\theta}$. By [27, Lem. 2.2.6] the input maps $\Phi_{t}\in\mathcal{B}(L^{2}(0,t;U),X)$ of the system $(A-\kappa BB^{\ast},B,B^{\ast})$, defined by $\Phi_{t}u=\int_{0}^{t}T_{\kappa}(t-s)Bu(s)\,\mathrm{d}s$ for $t\geq 0$ and $u\in L^{2}(0,t;U)$, have uniformly bounded operator norms; see also [16, Thm. 6.5.6], [31, Cor. 6.1]. Since $\|y(t)\|\lesssim\|x(t)\|^{\gamma}$ for all $t\geq t_{0}$, and since $\|x(t)\|=O(t^{-1/(2\beta)})$ as $t\to\infty$ by the first part of the result, we deduce that
@@
    The class of functions $\phi\colon U\to U$ to which the Theorem 2.1 can be applied includes radial functions defined by $\phi(0)=0$ and
    $\phi(u)=\psi(\|u\|)\frac{u}{\|u\|},\qquad u\in U\setminus\{0\},$
    where $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is a non-decreasing and locally Lipschitz continuous function satisfying $\psi(0)=0$. Then $\phi$ is monotone and locally Lipschitz continuous, and (1.2) holds provided there exists $\delta>0$ such that $\psi(r)\gtrsim r$ for $0\leq r\leq\delta$. Furthermore, if $\psi$ is twice continuously differentiable with $\psi^{\prime}(0)>0$ a simple estimate using Taylor’s theorem shows that (2.2) holds with $\kappa=\psi^{\prime}(0)$ for some $\varepsilon>0$ and some $\gamma\geq 2$. A concrete example is the function $\psi$ defined by $\psi(r)=\tanh(r)$ for $r\geq 0$. In this case, there exists $\varepsilon>0$ such that (2.2) holds for $\kappa=1$ and $\gamma=3$.
@@
    If in addition to the assumptions made in the first part of Theorem 2.1 we assume that $\phi$ is globally Lipschitz continuous and satisfies $\operatorname*{Re}\langle\phi(u),u\rangle\gtrsim\|u\|^{2}$ for all $u\in U$, then a slight adaptation of the proof yields uniform polynomial decay of the form
    $\|x(t)\|\lesssim\frac{\|x(0)\|_{D(A)}}{(1+t)^{1/(2\beta)}},\qquad t\geq 0,$
    for all classical solutions of (1.1), where $\|\cdot\|_{D(A)}$ denotes the graph norm of $A$. This conclusion is analogous to earlier results showing that exponential stability may be preserved under global growth conditions on the non-linear damping; see for instance [23].
As a simple consequence of Theorem 2.1 we obtain the following improvement of [11, Thm. 4.4] and [28, Thm. 4.2] in the linear case. Both of these earlier results include the assumption that $\beta\in(0,1]$, and in [11, Thm. 4.4] the further assumption that $D(A)=D(A^{*})$ is added. Our present approach requires neither. As before, we assume that $A$ is the generator of a contraction semigroup $(T(t))_{t\geq 0}$ on a Hilbert space $X$, and that $B\in\mathcal{B}(U,X)$ for some other Hilbert space $U$. Moreover, we let $A_{B}=A-BB^{*}$ with domain $D(A_{B})=D(A)$, and we let $(T_{B}(t))_{t\geq 0}$ denote the contraction semigroup on $X$ generated by $A_{B}$.
@@
Applying Theorem 2.1 with $\phi\colon U\to U$ taken to be the identity map we obtain $i\mathbb{R}\subseteq\rho(A_{B})$ and $\|T_{B}(t)x_{0}\|=O(t^{-1/(2\beta)})$ as $t\to\infty$ for all $x_{0}\in D(A)$. Since the semigroup $(T_{B}(t))_{t\geq 0}$ is contractive and $D(A)$ is dense in $X$ a simple approximation argument shows that $\|T_{B}(t)x_{0}\|\to 0$ as $t\to\infty$ for all $x_{0}\in X$. Moreover, an application of the uniform boundedness theorem yields $\|T_{B}(t)\smash{A_{B}^{-1}}\|=O(t^{-1/(2\beta)})$ as $t\to\infty$, and hence $\|T_{B}(t)x_{0}\|=o(t^{-1/(2\beta)})$ as $t\to\infty$ for all $x_{0}\in D(A)$ by [10, Thm. 2.4]. ∎
@@
Let us denote the eigenvalues of $A$ by $is_{k}$, $k\geq 1$, where $s_{k}\in\mathbb{R}$ for all $k$. By assumption there exists $\delta>0$ such that $|s_{k}-s_{j}|\geq\delta$ for all $j,k\geq 1$ with $j\neq k$. Let $\beta=\alpha/2$. By [11, Prop. 5.1] we have
@@
for all $k\geq 1$ and $x\in\operatorname*{Ker}(is_{k}-A)$. For $k\geq 1$, let $e_{k}$ be a normalised eigenvector corresponding to the eigenvalue $is_{k}$. Then $\{e_{k}:k\geq 1\}$ is an orthonormal basis for $X$, and it follows from (2.4) that $(1+|s_{k}|^{2})^{\beta}\|B^{\ast}e_{k}\|^{2}\gtrsim 1$ for all $k\geq 1$. Let $\tau>2\pi/\delta$. It follows from Ingham’s inequality [6, Prop. 1.5.2] that
@@
    The estimate in (2.4) is sometimes referred to as a wavepacket condition. The above proof shows that, under a uniform spectral gap condition, we may pass from a resolvent estimate for $A_{B}$ to a wavepacket condition and thence to an observability estimate. In [11, Sect. 3] it is shown, conversely, that even without the spectral gap condition general wavepacket conditions of this kind can be used to obtain resolvent estimates for $A_{B}$, which in turn imply decay rates for classical orbits of the damped semigroup $(T_{B}(t))_{t\geq 0}$.
@@
    By applying more sophisticated versions of Ingham’s inequality such as [6, Cor. 1.5.4] we may weaken the uniform gap condition to allow for a degree of repetition and clustering in the eigenvalues of $A$.
@@
Suppose that $A$ is skew-adjoint, $\sigma(A)$ consists of simple and uniformly separated eigenvalues, and that $i\mathbb{R}\subseteq\rho(A_{B})$ and $\|(is-A_{B})^{-1}\|\lesssim 1+|s|^{\alpha}$ for some $\alpha>0$ and all $s\in\mathbb{R}$. Suppose further that $\phi:U\to U$ is monotone and locally Lipschitz continuous and that (1.2) holds for some $\delta>0$. Then all generalised solutions of (1.1) satisfy $\|x(t)\|\to 0$ as $t\to\infty$, and all classical solutions of (1.1) satisfy $\|x(t)\|=O(t^{-1/\alpha})$ as $t\to\infty$.

### v2

@@ 1. Introduction
Motivated by applications to hyperbolic partial differential equations with non-linear damping, we consider abstractsemi-linear systems of the form
@@
to be solved subject to the initial condition $x(0)=x_{0}$. Here $A$ is assumed to be the generator of a contraction semigroup $(T(t))_{t\geq 0}$ on a Hilbert space $X$, $x_{0}\in X$ and $B$ is a bounded linear operator from another Hilbert space $U$ into $X$. Furthermore, $\phi\colon U\to U$ is a locally Lipschitz continuous map which satisfies $\phi(0)=0$ and is monotone in the sense that
@@
The study of polynomial rates of decay for non-linearly damped evolution equations goes back to the work of Haraux and others in the 1970s and 1980s. These early works consider wave equations and abstract second-order systems subject to non-linear viscous damping, establishing rates of energy decay typically by means of Lyapunov analysis and multiplier methods [22, 23, 24, 31, 32, 45]; we refer the reader to [46, Chap. 5 & 6] for an overview of several of these developments. More recent works on rational and generalised decay rates for semigroups and (abstract) wave equations with non-linear viscous-type damping include [2, 4, 13, 25, 33], and the case of non-linear boundary damping has been studied in [1, 27, 42, 43, 44]. Exponential stability in the presence of non-linear damping has been investigated for instance in [14, 20, 29, 41].
One typical feature of these models is that in the special case of linear damping one obtains uniform exponential stability of the underlying semigroup; it is the non-linearity of the damping that gives rise to polynomial (or slower) rates of decay. Our assumptions are fundamentally different. For one thing, in our case the linear system generally fails to be uniformly exponentially stable, and indeed we are interested in cases where the non-linearity preserves the polynomial rate of decay one obtains for (classical) solutions of the corresponding linear system. Furthermore, the conditions we impose on the non-linear function $\phi$ are stated in terms of the inner product on the Hilbert space $U$, whereas most previous works in the setting of wave-type equations consider viscous damping acting pointwise in space. As a result, we follow a rather different approach based on a non-uniform observability estimate. In doing so we extend to the non-linear setting various results in [8]; the same results have previously been extended to abstract wave-type equations with non-linear viscous damping in [3, 5, 6, 36]. Our results may be viewed as providing rates of decay in earlier unquantified studies on asymptotic stability with non-linear damping such as [15, 16, 26, 28, 37].
A function $x(\cdot)\in C^{1}(\mathbb{R}_{+};X)$ is said to be a classical solution of (1.1) if $x(t)\in D(A)$ for all $t\geq 0$ and (1.1) holds for all $t\geq 0$. A function $x(\cdot)\in C(\mathbb{R}_{+};X)$ is said to be a mild solution of (1.1) if
$x(t)=T(t)x_{0}-\int_{0}^{t}T(t-s)B\phi(B^{*}x(s))\,\mathrm{d}s,\qquad t\geq 0.$
Note that if $\phi$ is a linear map then the notions of classical and mild solutions are consistent with the terminology used for linear abstract Cauchy problems. We recall, moreover, that in the linear case the mild solution is given by the semigroup orbit, and that it is a classical solution precisely when $x_{0}\in D(A)$.
@@
and in particular monotonicity of $\phi$ implies that any classical solution has non-increasing norm. It follows from [17, Thm. 11.1.5(b)] that (1.1) has a unique classical solution whenever $x_{0}\in D(A)$, and by density of $D(A)$ in $X$ there exists a unique mild solution for all $x_{0}\in X$. In fact, by monotonicity of $\phi$ any two solutions $x_{1}(\cdot),x_{2}(\cdot)$ satisfy
@@
and for any classical solution $x(\cdot)$ the function $\|\dot{x}(\cdot)\|$ is non-increasing by [30, Cor. 3.7].
@@ 2. Polynomial stability of abstract semi-linear systems
Suppose that $\phi\colon U\to U$ is monotone, locally Lipschitz continuous and satisfies $\phi(0)=0$, and that (1.2) holds for some $\delta>0$. Suppose further that there exist $\beta,\tau,c_{\tau}>0$ such that
@@
Then $i\mathbb{R}\subseteq\rho(A_{B})$. Furthermore, all mild solutions of (1.1) satisfy $\|x(t)\|\to 0$ as $t\to\infty$, and all classical solutions of (1.1) satisfy $\|x(t)\|=O(t^{-1/(2\beta)})$ as $t\to\infty$.
@@
where the implicit constants are independent of $k$. Noting that $x(k\tau)\in D(A)$, we may apply the moment inequality [18, Thm. II.5.34] to obtain
@@
for all sufficiently large $k\geq 1$ by [7, Lem. 1.3.4]. By monotonicity of $\|x(\cdot)\|$ we deduce that
@@
for all sufficiently large $t\geq 0$. Next we observe that, for any $\delta\geq\|B^{*}\|\|x_{0}\|$, the identity function on $U$ defined by $\phi(u)=u$ for all $u\in U$ satisfies our assumptions with $t_{0}=0$ and $L_{\delta}=1$. We deduce that $\|T_{B}(t)(I-A_{B})^{-1}\|=O(t^{-1/(2\beta)})$ as $t\to\infty$, where $(T_{B}(t))_{t\geq 0}$ is the $C_{0}$-semigroup of contractions generated by $A_{B}$. It follows from [10, Thm. 1.1] that $i\mathbb{R}\subseteq\rho(A_{B})$. Moreover, $(T_{B}(t))_{t\geq 0}$ is strongly stable in the sense that $\|T_{B}(t)x\|\to 0$ as $t\to\infty$ for all $x\in X$, by a standard density argument. It then follows from [16, Thm. 2.2] (noting that the result carries over, with the appropriate modifications, to the setting of complex Hilbert spaces, and that the assumptions of compact resolvent and approximate observability can be replaced by strong stability of $(T_{B}(t))_{t\geq 0}$) that whenever $\phi$ satisfies the conditions of our theorem we have $\|x(t)\|\to 0$ as $t\to\infty$ for all mild solutions of (1.1). Hence the first part of the proof shows that $\|x(t)\|=O(t^{-1/(2\beta)})$ as $t\to\infty$ for all classical solutions of (1.1), which completes the proof of the first part of the result.
Now assume, in addition, that $\|(is-A_{B})^{-1}\|\lesssim 1+|s|^{\alpha}$ for some $\alpha>0$ and all $s\in\mathbb{R}$, and that there exist $\gamma>\alpha/2+1$ and $\kappa,\varepsilon>0$ such that (2.2) holds. If $\alpha\geq 2\beta$ then the result already follows without any of the additional assumptions from what has already been proved, so we may assume that $\alpha<2\beta$. Recall from the first part of the proof that $\|x(t)\|\to 0$ as $t\to\infty$ for all mild solutions of (1.1), and that $\|x(t)\|=O(t^{-1/(2\beta)})$ as $t\to\infty$ for all classical solutions of (1.1). Note also that by [12, Lem. 2.11(c)] the operator $A_{\kappa}=A-\kappa BB^{\ast}$ satisfies $i\mathbb{R}\subseteq\rho(A_{\kappa})$ and $\|T_{\kappa}(t)A_{\kappa}^{-1}\|=O(t^{-1/\alpha})$ as $t\to\infty$, where $(T_{\kappa}(t))_{t\geq 0}$ is the contraction semigroup generated by $A_{\kappa}$. Let $x(\cdot)$ be a classical solution of (1.1), and let $t_{0}\geq 0$ be such that $\|B^{*}x(t)\|\leq\varepsilon$ for all $t\geq t_{0}$. We have
@@
for all $t\geq t_{0}$. Since $x(t)\in D(A)$ for all $t\geq 0$, we have $\|T_{\kappa}(t-t_{0})x(t_{0})\|=O(t^{-1/\alpha})$ as $t\to\infty$. Moreover, since $\phi$ is locally Lipschitz continuous and $Bx(\cdot)$ is continuously differentiable, we have $\int_{t_{0}}^{t_{\theta}}T_{\kappa}(t_{\theta}-s)By(s)\,\mathrm{d}s\in D(A)$ for all $t\geq t_{0}$ by [9, Cor. 3.1.17]. The second term in (2.3) therefore satisfies
@@
where $y_{\theta}(t)=0$ for $t\in[0,t_{\theta})$ and $y_{\theta}(t)=y(t)$ for $t\geq t_{\theta}$. By [34, Lem. 2.2.6] the maps $\Phi_{t}\in\mathcal{B}(L^{2}(0,t;U),X)$, defined by $\Phi_{t}u=\int_{0}^{t}T_{\kappa}(t-s)Bu(s)\,\mathrm{d}s$ for $t\geq 0$ and $u\in L^{2}(0,t;U)$, have uniformly bounded operator norms; see also [17, Thm. 6.5.6], [38, Cor. 6.1]. Since $\|y(t)\|\lesssim\|x(t)\|^{\gamma}$ for all $t\geq t_{0}$, and since $\|x(t)\|=O(t^{-1/(2\beta)})$ as $t\to\infty$ by the first part of the result, we deduce that
@@
    The class of functions $\phi\colon U\to U$ to which the Theorem 2.1 can be applied includes radial functions defined by $\phi(0)=0$ and ϕ(u)=ψ(∥u∥)u∥u∥,  u∈U∖{0}, where $\psi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ is a non-decreasing and locally Lipschitz continuous function satisfying $\psi(0)=0$. Then $\phi$ is monotone and locally Lipschitz continuous, and (1.2) holds provided there exists $\delta>0$ such that $\psi(r)\gtrsim r$ for $0\leq r\leq\delta$. Furthermore, if $\psi$ is twice continuously differentiable with $\psi^{\prime}(0)>0$ a simple estimate using Taylor’s theorem shows that (2.2) holds with $\kappa=\psi^{\prime}(0)$ for some $\varepsilon>0$ and some $\gamma\geq 2$. A concrete example is the function $\psi$ defined by $\psi(r)=\tanh(r)$ for $r\geq 0$. In this case, there exists $\varepsilon>0$ such that (2.2) holds for $\kappa=1$ and $\gamma=3$.
@@
    If in addition to the assumptions made in the first part of Theorem 2.1 we assume that $\phi$ is globally Lipschitz continuous and satisfies $\operatorname*{Re}\langle\phi(u),u\rangle\gtrsim\|u\|^{2}$ for all $u\in U$, then a slight adaptation of the proof yields uniform polynomial decay of the form ∥x(t)∥ ≲∥x(0)∥D(A)(1+t)1/(2β) ,   t≥0, for all classical solutions of (1.1), where $\|\cdot\|_{D(A)}$ denotes the graph norm of $A$. This conclusion is analogous to earlier results showing that exponential stability may be preserved under global growth conditions on the non-linear damping; see for instance [27].
  4. (d)
    It is clear that the above proof relies crucially on boundedness of $B$. On the other hand, the results in [1, 8, 12] allow for certain ‘unbounded’ operators $B$ mapping into an extrapolation space of $X$. We leave open to what extent our results can be generalised to this considerably more challenging setting.
As a simple consequence of Theorem 2.1 we obtain the following improvement of [12, Thm. 4.4] and [35, Thm. 4.2] in the linear case. Both of these earlier results include the assumption that $\beta\in(0,1]$, and in [12, Thm. 4.4] the further assumption that $D(A)=D(A^{*})$ is added. Our present approach requires neither. As before, we assume that $A$ is the generator of a contraction semigroup $(T(t))_{t\geq 0}$ on a Hilbert space $X$, and that $B\in\mathcal{B}(U,X)$ for some other Hilbert space $U$. Moreover, we let $A_{B}=A-BB^{*}$ with domain $D(A_{B})=D(A)$, and we let $(T_{B}(t))_{t\geq 0}$ denote the contraction semigroup on $X$ generated by $A_{B}$.
@@
Applying Theorem 2.1 with $\phi\colon U\to U$ taken to be the identity map we obtain $i\mathbb{R}\subseteq\rho(A_{B})$ and $\|T_{B}(t)x_{0}\|=O(t^{-1/(2\beta)})$ as $t\to\infty$ for all $x_{0}\in D(A)$. Since the semigroup $(T_{B}(t))_{t\geq 0}$ is contractive and $D(A)$ is dense in $X$ a simple approximation argument shows that $\|T_{B}(t)x_{0}\|\to 0$ as $t\to\infty$ for all $x_{0}\in X$. Moreover, an application of the uniform boundedness theorem yields $\|T_{B}(t)\smash{A_{B}^{-1}}\|=O(t^{-1/(2\beta)})$ as $t\to\infty$, and hence $\|T_{B}(t)x_{0}\|=o(t^{-1/(2\beta)})$ as $t\to\infty$ for all $x_{0}\in D(A)$ by [11, Thm. 2.4]. ∎
@@
Let us denote the eigenvalues of $A$ by $is_{k}$, $k\geq 1$, where $s_{k}\in\mathbb{R}$ for all $k$. By assumption there exists $\delta>0$ such that $|s_{k}-s_{j}|\geq\delta$ for all $j,k\geq 1$ with $j\neq k$. Let $\beta=\alpha/2$. By [12, Prop. 5.1] we have
@@
for all $k\geq 1$ and $x\in\operatorname*{Ker}(is_{k}-A)$. For $k\geq 1$, let $e_{k}$ be a normalised eigenvector corresponding to the eigenvalue $is_{k}$. Then $\{e_{k}:k\geq 1\}$ is an orthonormal basis for $X$, and it follows from (2.4) that $(1+|s_{k}|^{2})^{\beta}\|B^{\ast}e_{k}\|^{2}\gtrsim 1$ for all $k\geq 1$. Let $\tau>2\pi/\delta$. It follows from Ingham’s inequality [7, Prop. 1.5.2] that
@@
    The estimate in (2.4) is sometimes referred to as a wavepacket condition. The above proof shows that, under a uniform spectral gap condition, we may pass from a resolvent estimate for $A_{B}$ to a wavepacket condition and thence to an observability estimate. In [12, Sect. 3] it is shown, conversely, that even without the spectral gap condition general wavepacket conditions of this kind can be used to obtain resolvent estimates for $A_{B}$, which in turn imply decay rates for classical orbits of the damped semigroup $(T_{B}(t))_{t\geq 0}$.
@@
    By applying more sophisticated versions of Ingham’s inequality such as [7, Cor. 1.5.4] we may weaken the uniform gap condition to allow for a degree of repetition and clustering in the eigenvalues of $A$.
@@
Suppose that $A$ is skew-adjoint, $\sigma(A)$ consists of simple and uniformly separated eigenvalues, and that $i\mathbb{R}\subseteq\rho(A_{B})$ and $\|(is-A_{B})^{-1}\|\lesssim 1+|s|^{\alpha}$ for some $\alpha>0$ and all $s\in\mathbb{R}$. Suppose further that $\phi\colon U\to U$ is monotone and locally Lipschitz continuous and that (1.2) holds for some $\delta>0$. Then all mild solutions of (1.1) satisfy $\|x(t)\|\to 0$ as $t\to\infty$, and all classical solutions of (1.1) satisfy $\|x(t)\|=O(t^{-1/\alpha})$ as $t\to\infty$.